# Ayoung Kim, Dept. of Mechanical Engineering, SNU, Seoul, S. Korea

### v3

@@ Abstract
Maritime environmental sensing requires overcoming challenges from complex conditions such as harsh weather, platform perturbations, large dynamic objects, and the requirement for long detection ranges. While cameras and LiDAR are commonly used in ground vehicle navigation, their applicability in maritime settings is limited by range constraints and hardware maintenance issues. Radar sensors, however, offer robust long-range detection capabilities and resilience to physical contamination from weather and saline conditions, making it a powerful sensor for maritime navigation. Among various radar types, X-band radar (e.g., marine radar) is widely employed for maritime vessel navigation, providing effective long-range detection essential for situational awareness and collision avoidance. Nevertheless, it exhibits limitations during berthing operations where close-range object detection is critical. To address this shortcoming, we incorporate W-band radar (e.g., Navtech imaging radar), which excels in detecting nearby objects with a higher update rate. We present a comprehensive maritime sensor dataset featuring multi-range detection capabilities. This dataset integrates short-range LiDAR data, medium-range W-band radar data, and long-range X-band radar data into a unified framework. Additionally, it includes object labels for oceanic object detection usage, derived from radar and stereo camera images. The dataset comprises seven sequences collected from diverse regions with varying levels of estimation difficulty, ranging from easy to challenging, and includes common locations suitable for global localization tasks. This dataset serves as a valuable resource for advancing research in place recognition, odometry estimation, SLAM (SLAM), object detection, and dynamic object elimination within maritime environments. Dataset can be found in following link: https://sites.google.com/view/rpmmoana
@@ 1 Introduction
The existing maritime dataset (Chung et al., 2023) represents a significant contribution, providing navigational data from marine radar only with X-band wavelength, along with LiDAR, and camera systems and ground truth information. While this dataset marks a pioneering step for maritime research, it lacks regional diversity, limiting its applicability across varied navigational environments. Understanding this limitation requires familiarity with the stages of vessel navigation: berthing at a port, sailing in open waters, and docking. While near-range detection is critical during berthing and docking, long-range detection is essential for open-water navigation. Existing datasets provide limited coverage of such wide-open maritime areas, restricting utility for radar navigation research.
@@
Moreover, the existing dataset faces limitations in sensor capabilities, as the canal environment is insufficient to represent radar-based marine navigation. Traditional marine radars (X-band) are pivotal in ensuring situational awareness and preventing collisions in wide oceanic areas with their long-range detection ability. Despite its broad sensing range, X-band radars are limited in detecting short-range objects and often suffer from multipath interference and noise in the canal area, as illustrated in Figure. 1. Therefore, it is necessary to explore alternative sensors for short-range detection. While LiDAR has proven effective for this short-range detection, it is not a sustainable solution in maritime environments due to challenges like sea fog, corrosion, and the necessity for frequent cleaning during extended voyages. We have found that W-band radar can be a superior alternative to LiDAR for short-range and wide-area detection in marine environments. W-band radar, often called scanning radar, has previously demonstrated its potential exclusively in ground vehicle applications due to its extended detection range compared to LiDAR. Its higher resolution and environmentally robust detection ability can mitigate the shortcomings of existing sensors in the maritime domain. Nevertheless, there have been no reports of its use in marine environments.
@@
    The MOANA dataset represents the first multi-radar (X-band and W-band) dataset collected in a maritime environment. The X-band radar is utilized for wide-area detection, while the W-band radar delivers high-resolution imaging for near-port areas. Their differing bandwidths allow for seamless integration, and our benchmark results also demonstrate that a complementary sensor configuration enables robust navigation across various maritime conditions. Our dataset also provides cameras and LiDAR data for the ground truth labels of detected objects, encouraging diverse fusion strategies for vessel navigation tasks: berthing, sailing, and docking.
@@
    The MOANA dataset tackles maritime navigation tasks in varying environments. Our dataset encompasses both structured port and unstructured ocean and island settings with varying vessel sizes. This large spectrum of scene diversity underscores the necessity of navigation algorithms operating in multifarious environments. Additionally, the loops and overlaps between sequences support inter and intra-place recognition in maritime environments.
@@ 2 Related works
This section summarizes the existing maritime dataset focused on cameras and LiDAR and examines W-band datasets that have so far been limited to ground-based usage in advance of introducing the first multi-radar dataset to incorporate the W-band radar from maritime environments. The brief summarization is depicted in Table. 1.
@@ 2.1 Marine Robotics Datasets
The development of maritime environment datasets also began with vision-based approaches. Early efforts included a camera-infrared dataset for day and low-light conditions (Zhang et al., 2015) and a bird’s-eye-view maritime surveillance camera dataset (Ribeiro et al., 2017). Additionally, maritime imagery for learning-based models has been provided (Bovcon et al., 2019). Attempts to integrate camera, LiDAR, and single-chip radar on unmanned surface vehicles were made (Cheng et al., 2021), but the single-chip radar system has insufficient resolution for oceanic environments. Recently, (Chung et al., 2023) introduced a multi-purpose marine radar dataset. While camera and LiDAR data exhibit limitations in open water scenarios due to detection range constraints, including X-band marine radar in this dataset addresses these shortcomings, offering a more comprehensive solution for maritime applications. Unlike existing maritime datasets, the MOANA dataset incorporates both X-band and W-band radar to enhance the robustness of detectability for diverse maritime navigation tasks.
@@ 2.2 W-band Radar Datasets
The use of W-band radar has been confined to ground vehicle datasets. Pioneering radar datasets, including MulRan (Kim et al., 2020) and the Oxford Radar Robotcar (Barnes et al., 2020), have established radar as a viable sensor for navigation. Following these foundational efforts, datasets have been extended to include various challenging environments, such as adverse weather conditions (Sheeny et al., 2021), multi-season driving (Burnett et al., 2023), and complex terrains (Gadd et al., 2024). The MOANA dataset is differentiated from the existing W-band radar datasets, exploiting W-band radar in the maritime domain for the first time. Consequently, the MOANA dataset serves as a milestone to verify the potential of the W-band radar in vessel navigation research.
@@ 3.2.1 Multiple Radar Calibration
Two radar data are provided as image types that are discriminated against with image pixels. We convert the W-band polar coordinate images to Cartesian coordinates with the same resolution as the X-band radar and perform the extrinsic calibration by considering the overlap of the prominent features at the pixel level. Since the two radars are structurally incapable of providing information in the vertical direction, vertical extrinsic calibration was conducted using a CAD model.
@@ 3.2.2 LiDAR W-band Radar Camera Calibration
Table 2: The radar specifications utilized in dataset
Dataset Manufacture Model Range Resolution Range
Port,Island SIMRAD HALO4 3m 1852m
Port Navtech RAS6-DEV-A 0.175m 600m
Island Navtech RAS6-DEV-X 0.175m 600m
@@ 4.1 Data Composition
As detailed in Chap. 3.1, our dataset includes data from two different band radars, LiDAR, stereo camera, and GNSS Reciever. Figure. 3 provides the data sample with detection range comparison for each sensor. Both scanning radars and stereo cameras generate image data, making this dataset composed of four images for every data publication. To address potential challenges in managing large data volumes, we offer an individual file structure in $Sequence/sensor\_data/sensor\_type$ directory, enabling users to download only the data relevant to their needs. Additionally, we provide a ROS-based data publisher to facilitate seamless access and integration of the dataset. An overview of the file structure is illustrated in Figure. 4.
Figure 4: Data organization and file structure for each sequence of MOANA dataset. Ground truth poses are provided as text files. Extrinsic calibration parameters for the camera, LiDAR, and radars are defined relative to the base frame. Sensor data is available with distinguished file formats, such as PNG images. Additionally, labels for each frame are supplied as JSON files, each named according to the corresponding frame timestamp. For the stereo camera annotation, we denote the left camera as Cam0, and the right camera as Cam1 for JSON files.
@@ 4.1.1 X-Band Radar
X-band radar data, commonly referred to as marine radar data, is stored as Cartesian coordinate image files. The radar’s maximum range is set to 1 nautical mile for the Island sequences and 1 km for the Port sequences. To optimize sensor fusion efficiency, we fixed the sensor update rate at 1 Hz. Originally, the radar requires three seconds to complete a full 360-degree rotation. However, the data provided updates one-third of the full image per second. This modification was implemented to reduce dynamic changes that occur over short time intervals, thereby enhancing the temporal resolution of the sensor data.
@@ 4.1.2 W-Band Radar
W-band radar, commonly referred to as imaging radar, was employed in our study using the Navtech RAS6 models. This radar system has a range of approximately 600 meters and provides data in the form of polar coordinate images. Two variants of the W-band radar were utilized for different experimental sequences. For the Port sequences, we used the RAS6-DEV-A model, characterized by radar rays that propagate horizontally, detecting objects at the same elevation level as the radar itself. In contrast, the Island sequences utilized the RAS6-DEV-X model, which emits rays capable of detecting objects below the radar’s elevation. This adaptation was necessary to accommodate the vessel’s size differences; specifically, the vessel in the Island sequences has a height of approximately 6 meters, double that of the vessel used in the Port sequences. Although W-band radar is traditionally utilized for ground vehicle navigation, it proves to be highly advantageous for maritime applications. It offers superior resolution in detecting surrounding environments compared to X-band radar and exhibits greater robustness than sensors based on optical wavelengths. The specification of the radar configuration is detailed in Table. 2.
@@ 4.1.3 LiDAR and Camera
To validate the environmental conditions during sensor data acquisition, we additionally equipped the vessel with fundamental navigational sensors, LiDAR and a stereo camera. Due to LiDAR’s limited detection range, most of its data is sparse except in the berthing regions. The primary reason for incorporating LiDAR into this dataset is to assess precision during berthing maneuvers, where radar saturation can occur.
The stereo camera data allowed us to detect nearby vessels and structures even when operating in open ocean. This visual data serves as ground truth for object labeling tasks. Estimating visual odometry using only the stereo camera is challenging in oceanic environments due to the scarcity of discernible objects; however, this presents an additional challenge for researchers to address.
@@ 4.1.4 Calibration and Ground Truth
The extrinsic calibration files and ground truth positions from the GNSS receiver are provided as text files. We employed the Hemisphere V500 GNSS receiver, which delivers precise positioning along with heading information. While the receiver’s native update frequency is 10 Hz, we have synchronized the data with other sensors to a 1 Hz update rate in our dataset. The GNSS receiver serves as the reference frame for the vessel model, and the extrinsic parameters are calculated based on the correlation with this base frame.
@@ 4.1.5 Annotation Label
We provide the ground truth of 2D bounding box annotations for detected objects in X-band radar, W-band radar, and stereo images in the Single Island sequence. The annotations in the JSON file contain [$timestamp$, $id$, $category\_id$, $xmin$, $ymin$, $width$, $height$]. $id$ represents the tracked object ID that is detected in consecutive frames. The $category\_id$ is fixed at a value of 1, corresponding to the objects that should be avoided during navigation (e.g. vessels or buoys). The [$xmin$, $ymin$, $width$, $height$] parameters represent the bounding box location and dimensions of the detected objects in each image. For the W-band radar, the locations of the bounding box are represented in 1024$\times$1024 Cartesian image coordinates, which are converted from the polar coordinate image. Figure. 5 depicts an example of provided 2D annotations in each images.

### v4

@@ Abstract
Maritime environmental sensing requires overcoming challenges from complex conditions such as harsh weather, platform perturbations, large dynamic objects, and the requirement for long detection ranges. While cameras and LiDAR are commonly used in ground vehicle navigation, their applicability in maritime settings is limited by range constraints and hardware maintenance issues. Radar sensors, however, offer robust long-range detection capabilities and resilience to physical contamination from weather and saline conditions, making it a powerful sensor for maritime navigation. Among various radar types, X-band radar (e.g., marine radar) is widely employed for maritime vessel navigation, providing effective long-range detection essential for situational awareness and collision avoidance. Nevertheless, it exhibits limitations during berthing operations where near-field detection is critical. To address this shortcoming, we incorporate W-band radar (e.g., Navtech imaging radar), which excels in detecting nearby objects with a higher update rate. We present a comprehensive maritime sensor dataset featuring multi-range detection capabilities. This dataset integrates short-range LiDAR data, medium-range W-band radar data, and long-range X-band radar data into a unified framework. Additionally, it includes object labels for oceanic object detection usage, derived from radar and stereo camera images. The dataset comprises seven sequences collected from diverse regions with varying levels of navigation algorithm estimation difficulty, ranging from easy to challenging, and includes common locations suitable for global localization tasks. This dataset serves as a valuable resource for advancing research in place recognition, odometry estimation, SLAM, object detection, and dynamic object elimination within maritime environments. Dataset can be found in following link: https://sites.google.com/view/rpmmoana
@@ 1 Introduction
The existing maritime dataset (Chung et al., 2023) represents a significant contribution, providing navigational data from marine radar only with X-band wavelength, along with LiDAR (LiDAR), camera systems, and ground truth information. While this dataset marks a pioneering step for maritime research, it lacks regional diversity, limiting its applicability across varied navigational environments. Understanding this limitation requires familiarity with the stages of vessel navigation: berthing at a port, sailing in open waters, and docking. While near-range detection is critical during berthing and docking, long-range detection is essential for open-water navigation. Existing datasets provide limited coverage of such wide-open maritime areas, restricting utility for radar navigation research.
@@
Moreover, the existing dataset faces limitations in sensor capabilities, as the canal environment is the narrow area generating multipath effects in radar imagery, which is insufficient to represent radar-based navigation. Traditional marine radars (X-band) are pivotal in ensuring situational awareness and preventing collisions in wide oceanic areas with their long-range detection ability. Despite its broad sensing range, X-band radars are limited in detecting small objects near the vessels and often suffer from multipath interference and noise, as illustrated in Figure. 1. Therefore, it is necessary to explore alternative sensors for these types of detection. While LiDAR has proven its effectiveness for the high-resolution perception, it is not a sustainable solution in maritime environments due to challenges like sea fog, corrosion, and the necessity for frequent cleaning during extended voyages. We have found that W-band radar can be a superior alternative to LiDAR for short-range detection in marine environments. W-band radar, often called scanning radar, has previously demonstrated its potential exclusively in ground vehicle applications (Harlow et al., 2024) due to its extended detection range and robustness to the surroundings compared to LiDAR. Its environmentally resilient detection capabilities can effectively address the shortcomings of LiDAR in the maritime domain. Also, its higher resolution relative to the X-band radar provides comprehensive perception ability, which can be a reasonable alternative for LiDAR in short-range detection with high fidelity. Nevertheless, there have been no reports of its use in marine environments.
@@
    The MOANA dataset represents the first multi-radar (X-band and W-band) dataset collected in a maritime environment. The X-band radar is utilized for wide-area detection, while the W-band radar delivers high-resolution imaging for near-port areas. Their differing bandwidths allow for seamless integration, and our benchmark results demonstrate that a complementary sensor configuration enables robust navigation across various maritime conditions. Our dataset also integrates cameras and LiDAR data, encouraging diverse sensor fusion strategies for vessel navigation tasks: berthing, sailing, and docking. Moreover, the ground truth labels of detected objects for each modality are also provided for facilitating rigorous validation and benchmarking of multi-modal object detection in oceanic areas.
@@
    The MOANA dataset tackles maritime navigation tasks in varying environments. Our dataset encompasses both structured port and unstructured ocean and island settings with varying vessel sizes. This large spectrum of scene diversity underscores the necessity of navigation algorithms operating in multifarious environments. Additionally, the loops and overlaps between sequences support inter- and intra-place recognition in maritime environments.
@@ 2 Related works
Prior to introducing our first multi-radar dataset MOANA, we summarize existing maritime datasets—primarily focused on cameras and LiDAR—as well as W-band radar datasets, which have been limited to ground vehicle applications. A summary is provided in Table 1.
@@ 2.1 Marine Robotics Datasets
The development of maritime environment datasets also began with vision-based approaches for object detection and segmentation tasks. Early efforts included a camera-infrared dataset for day and low-light conditions (Zhang et al., 2015) and a bird’s-eye-view maritime surveillance camera dataset (Ribeiro et al., 2017). Additionally, MaSTr1325 (Bovcon et al., 2019) provided maritime imagery for learning-based models. Extending beyond the image domain, multimodal datasets for sensor fusion have been presented. FloW (Cheng et al., 2021b) utilized a single-chip radar and cameras for object detection in the waterway. (Lin et al., 2022) exploited LiDAR and AIS (AIS) for detection and semantic mapping in the maritime environment.
Recently, datasets for maritime navigation have been presented following the increasing demand in this field. USVInland (Cheng et al., 2021a) attempted to integrate camera, LiDAR, and single-chip radar on unmanned surface vehicles, but the single-chip radar system has insufficient resolution for navigation in oceanic environments. RoboWhaler (DeFilippo et al., 2021) equipped INS (INS), LiDAR with marine radar in ASV (ASV) for navigation applications. MassMIND (Nirgudkar et al., 2023) integrated the long wave infrared camera on the RoboWhaler system to complement the lack of diversity of optical images in maritime environments. The Pohang Canal dataset (Chung et al., 2023) introduced a multi-purpose marine radar dataset. While camera and LiDAR data exhibit limitations in open water scenarios due to detection range constraints, including X-band marine radar in this dataset addresses these shortcomings, offering a more comprehensive solution for maritime applications. PoLaRIS (Choi et al., 2024) extended the Pohang Canal dataset, providing the object bounding box and tracking annotations. Unlike existing maritime datasets, the MOANA dataset incorporates both X-band and W-band radar to enhance the robustness of detectability for diverse maritime navigation tasks.
@@ 2.2 W-band Radar Datasets
The use of W-band radar has been confined to ground vehicle datasets. Pioneering W-band radar datasets, including the Oxford Radar RobotCar (Barnes et al., 2020) and MulRan (Kim et al., 2020) datasets, have established radar as a viable sensor for navigation. The Oxford Radar RobotCar dataset provides extensive radar data alongside complementary sensor modalities for robust localization and mapping. However, its focus on a single urban area limits environmental diversity. Similarly, the MulRan dataset offers valuable multi-modal data for urban place recognition but remains constrained by its emphasis on structured urban environments and a lack of adverse weather scenarios.
Following these foundational efforts, Radiate (Sheeny et al., 2021) incorporates adverse weather conditions, such as rain and snow, along with ground truth object labels, thereby demonstrating the weather-resilient perception capability of W-band radar in harsh conditions. The Boreas dataset (Burnett et al., 2023) further expands on this by encompassing multi-seasonal data with annotation labels for object detection, showcasing radar performance under diverse seasonal conditions. Together, the Radiate and Boreas datasets highlight the potential of W-band radar for robust perception in autonomous driving applications.
However, despite these advancements, the majority of existing W-band radar datasets remain focused on terrestrial, structured environments. Addressing this gap, OORD (Gadd et al., 2024) extends W-band radar usage from structured settings to off-road terrains, demonstrating its applicability in unstructured open environments. Recently, HeRCULES (Kim et al., 2025) exploits W-band radar with 4D automotive radar to encompass diverse weather conditions in congested environments with high dynamics.
The MOANA dataset is differentiated from these efforts by exploiting W-band radar in the maritime domain for the first time, thereby serving as a milestone for robust navigation in marine environments.
@@ 3.2.1 Multiple Radar Calibration
The X-band radar data is provided in the Cartesian coordinate image, whereas the W-band radar data is presented in the polar coordinate image. We convert the W-band polar coordinate images to Cartesian coordinates with the same resolution as the X-band radar and perform the extrinsic calibration by considering the overlap of the prominent features at the pixel level. Since the two radars are structurally incapable of providing information in the vertical direction, vertical extrinsic calibration was conducted using a CAD model.
@@ 4.1 Data Composition
As detailed in Section §3.1, our dataset includes data from two different band radars, LiDAR, stereo camera, and GNSS Reciever. Figure. 3 provides the data sample with detection range comparison for each sensor. Both scanning radars and stereo cameras generate image data, making this dataset composed of four images for every data publication. To address potential challenges in managing large data volumes, we offer an individual file structure in $Sequence/sensor\_data/sensor\_type$ directory, enabling users to download only the data relevant to their needs. Additionally, we provide a ROS-based data publisher to facilitate seamless access and integration of the dataset. An overview of the file structure is illustrated in Figure. 4.
Figure 4: Data organization and file structure for each sequence of MOANA dataset. GNSS poses are provided as text files. Extrinsic calibration parameters for the camera, LiDAR, and radars are defined relative to the base frame. Sensor data is available with distinguished file formats, such as PNG images. Additionally, labels for each frame are supplied as JSON files, each named according to the corresponding frame timestamp. For the stereo camera annotation, we denote the left camera as Cam0 and the right camera as Cam1 for JSON files.
(a) Satellite
(b) X-band Radar
(c) W-band radar
(d) LiDAR
Figure 5: Map constructions from GNSS data-based accumulation of the Island sequences. The X-band radar (b) provides full coverage of the dataset region due to its $km$ level range and strong penetration capability; however, it exhibits relatively low resolution and accuracy. In contrast, the W-band radar (c) offers high-resolution sensing but is limited to detecting only $600m$ ranges. In the berthing area (highlighted by the red-colored square), the X-band radar fails to generate a detailed map. Conversely, the W-band radar (c) and LiDAR (d) produce a high-resolution map, capturing fine structural details.
@@ 4.1.1 X-Band Radar
X-band radar data, commonly referred to as marine radar data, is stored as Cartesian coordinate image files in PNG format. The radar’s maximum range is set to 1 nautical miles (1,852 m) for the Port sequences and 1.5 nautical miles (2,778 m) for the Island sequences. Due to the large error bound in the X-band radar SDK, we manually refined the range estimates using satellite imagery, LiDAR, and W-band radar data. The adjusted ranges are 2,498 m for the Port sequence and 3,328 m for the Island sequence. We fixed the sensor update rate at 1 Hz to optimize sensor fusion efficiency. Originally, the radar required three seconds to complete a full 360-degree rotation. However, the data provided updates one-third of the full image per second. This modification was implemented to reduce dynamic changes that occur over short time intervals, thereby enhancing the temporal resolution of the sensor data.
@@ 4.1.2 W-Band Radar
W-band radar, commonly referred to as imaging radar, was employed in our dataset using the Navtech RAS6 models. This radar system has a range of approximately 600 meters and provides data in the form of polar coordinate images. Two variants of the W-band radar were utilized for different experimental sequences. For the Port sequences, we used the RAS6-DEV-A model, characterized by radar rays that propagate horizontally, detecting objects at the same elevation level as the radar itself. In contrast, the Island sequences utilized the RAS6-DEV-X model, which emits rays capable of detecting objects below the radar’s elevation. This adaptation was necessary to accommodate the vessel’s size differences; specifically, the vessel in the Island sequences has a height of approximately 6 meters, double that of the vessel used in the Port sequences. For consistency, the W-band data is saved in PNG format for both sequences. Although W-band radar is traditionally utilized for ground vehicle navigation, it proves to be highly advantageous for maritime applications. It offers superior resolution in detecting surrounding environments compared to X-band radar and exhibits greater robustness than sensors based on optical wavelengths. The detailed specification of the radar configuration is detailed in Table. 2.
Table 2: Radar Specifications (S: SIMRAD, N: Navtech)
Dataset Type Model Resolution Range Band Frequency
Port X-band HALO4 (S) 2.44 m/px 2498 m 9.410 $\sim$ 9.498 GHz
W-band RAS6-DEV-A (N) 0.175 m/px 600 m 76 $\sim$ 77 GHz
Island X-band HALO4 (S) 3.25 m/px 3328 m 9.410 $\sim$ 9.498 GHz
W-band RAS6-DEV-X (N) 0.175 m/px 600 m 76 $\sim$ 77 GHz
@@ 4.1.3 LiDAR and Camera
To validate the environmental conditions during sensor data acquisition, we additionally equipped the vessel with fundamental navigational sensors, LiDAR and a stereo camera. LiDAR data are stored in BIN files, containing the ($x$, $y$, $z$) coordinates and intensity index with float32 format. Due to LiDAR’s limited detection range, most of its data is sparse except in the berthing regions. The primary reason for incorporating LiDAR into this dataset is to assess precision during berthing maneuvers, where radar saturation can occur.
The stereo camera data, which are saved in JPG format, allowed us to detect nearby vessels and structures even when operating in the open ocean. This visual data serves as ground truth for object labeling tasks. Estimating visual odometry using only the stereo camera is challenging in oceanic environments due to the scarcity of discernible objects; however, this presents an additional challenge for researchers to address.
@@ 4.1.4 Calibration and Ground Truth
The extrinsic calibration files and positions from the GNSS receiver are provided as text files. We employed the Hemisphere V500 GNSS receiver, which delivers precise positioning along with heading information. The GNSS receiver serves as the reference frame for the vessel model, and the extrinsic parameters for other sensors are represented in this base frame. The GNSS-based maps are presented in Figure. 5. Both X-band and W-band radars produce accurate maps corresponding to their sensor specifications, and LiDAR generates detailed map in the narrow berthing region. However, signal reception failures occur in specific regions, leading to significant positional drift. To ensure rigorous odometry evaluation, a precisely adjusted ground truth is required. As part of future work, we aim to optimize all sensor data and generate a refined ground truth to facilitate high-precision state estimation using our dataset.
@@ 4.1.5 Annotation Label
We provide the ground truth of 2D bounding box annotations for detected objects in X-band radar, W-band radar, and stereo images in the Single Island sequence. The annotations in the JSON file contain [$timestamp$, $id$, $category\_id$, $xmin$, $ymin$, $width$, $height$]. $id$ represents the tracked object ID that is detected in consecutive frames. The $category\_id$ is fixed at a value of 1, corresponding to the objects that should be avoided during navigation (e.g. vessels or buoys). The [$xmin$, $ymin$, $width$, $height$] parameters represent the bounding box location and dimensions of the detected objects in each image. In X-band radar data, multipath effects often result in the presence of ghost objects or terrain artifacts. To ensure reliability, we annotated only genuine objects by cross-referencing information from additional sensor modalities. The provided bounding boxes assist in identifying potential false detections within the radar data. Furthermore, occluded objects tend to merge into a single cloud. To assign a track ID to these combined vessels, we employed a weighting scheme that considers vessel size, detection frequency, and spatial consistency. Finally, due to the limited resolution of X-band radar, certain vessels may appear fused with surrounding terrain. In such cases, vessel bounding boxes were delineated separately only when their contours could be distinctly identified. For the W-band radar, the locations of the bounding box are represented in 1024$\times$1024 Cartesian image coordinates, which are converted from the polar coordinate image. Figure. 6 depicts an example of provided 2D annotations in each images.